OpenStudy (anonymous):

solve the equation, 2y^2 = 5y confuse with this one.

sam (.sam.):

2y^(2)=5y 2y^(2)-5y=0 Factor out the GCF of y from 2y^(2)-5y. y(2y-5)=0 Set the single term factor on the left-hand side of the equation equal to 0. y=0 Set each of the factors of the left-hand side of the equation equal to 0. 2y-5=0 Since -5 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 5 to both sides. 2y=5 Divide each term in the equation by 2. y=(5)/(2) The complete solution is the set of the individual solutions. y=0,(5)/(2)

OpenStudy (callisto):

2y^2 = 5y 2y^2 -5y= 0 y(2y-5) =0 y=0 or 2y-5 =0 y=0 or y=5/2